# OpenScan: A Benchmark for Generalized Open-Vocabulary 3D Scene Understanding

### v2

@@ Appendix F Broader Impact
| Affordance | carry things | [object] is used for carrying things | holding up a roof | [object] is used for holding up a roof |
@@
| Property | useful for camping | [object] is useful for camping | soft | [object] is soft |
@@
| Type | baggage | [object] is a baggage | seat | [object] is a seat |
@@
| Manner | pack | [object] is a way of packing | observe | [object] is a way of observing |
@@
| Synonym | weight | [object] is related to weight | news | [object] is related to news |
@@
| Requirement | sit down | sitting down requires [object] | be unpluged | [object] does not desire to be unpluged |
@@
| Element | water | [object] has water | news | [object] has news |
@@
| Material | wood | [object] is made of wood | fabric | [object] is made of fabric |

### v3

@@ G Broader Impact
Affordance |
carry things | [object] is used for carrying things | holding up a roof | [object] is used for holding up a roof |
@@
Property |
useful for camping | [object] is useful for camping | soft | [object] is soft |
@@
Type |
baggage | [object] is a baggage | seat | [object] is a seat |
@@
Manner |
pack | [object] is a way of packing | observe | [object] is a way of observing |
@@
Synonym |
weight | [object] is related to weight | news | [object] is related to news |
@@
Requirement |
sit down | sitting down requires [object] | be unpluged | [object] does not desire to be unpluged |
@@
Element |
water | [object] has water | news | [object] has news |
@@
Material |
wood | [object] is made of wood | fabric | [object] is made of fabric |